OpenStudy (anonymous):

hey I'm doing edexcel igcse's here's the problem : simplify fully - 5/2x-6 - x+2/xsquared - 4x + 3

OpenStudy (anonymous):

its not -5 btw, it's just 5

OpenStudy (anonymous):

\[5/(2x-6) - (x+2)/(x^2 -4x +3) \] Is this the original equation?

OpenStudy (valpey):

If you factor the denominators of both terms you will see they have something in common. Then you just multiply each term on the top and bottom by the parts they are missing. It looks like this:\[\frac{5}{2x-6}-\frac{x+2}{x^2-4x+3}\]\[\frac{5}{2(x-3)}-\frac{x+2}{(x-1)(x-3)}\] Now we multiply the first term by (x-1)/(x-1) and the second by 2/2:\[\frac{5}{2x-6}-\frac{x+2}{x^2-4x+3}\]\[\frac{(x-1)5}{2(x-1)(x-3)}-\frac{2(x+2)}{2(x-1)(x-3)}\] Now we have common denominators.

OpenStudy (anonymous):

HEY, so apparently the answer was 3/2(x-1) .. At my exam I wrote 1.5/(x-l) as they said 'simplify fully'.. will I get marks cut off for that?